OpenStudy (pphalke):

A fraternity is to elect a president and a treasurer from the group of 40 members. How many ways can those two officers be elected?

OpenStudy (thomasfire55):

well it would be 40*39 do you know why

OpenStudy (thomasfire55):

40*39=1560

OpenStudy (pphalke):

no idea why but the answer is right 1,560 ways

OpenStudy (thomasfire55):

do you want me to explain why?

OpenStudy (oops):

This has to do with factorials (!) and permutations There's a very good khanacademy video that explains this: https://www.khanacademy.org/math/probability/probability-and-combinatorics-topic/permutations/v/permutation-formula If you don't have time to watch that, just know that the permutation formula is \[\frac{ n! }{ (n-r)! }\]with n being, in your scenario, the amount of members to choose from for prez and treasurer, and r being the two positions needed to be filled. \[\frac{ 40! }{ (40-2)! }\]